# Quantum property testing in sparse directed graphs

### v1

@@ 1.3 Our results and techniques
We will prove this theorem in Section 5. The authors in [bogdanov2002lower] prove classical hardness by using the hardness of distinguishing $k$-wise independent functions from uniform ones. By the polynomial method (again), this problem is also hard for quantum algorithms, and so we find a similar lower bound. In [yoshida2010query], the authors used various reductions to argue that a number of other problems are maximally hard, including Hamiltonian Path/Cycle, approximating Independent Set/Vertex Cover size etc. As a corollary of our quantum lower bound, we also obtain maximal quantum query complexity for these problems.
1.4 Related works on collision finding
The problem of collision finding is a ubiquitous problem in the field of algorithm theory with wide applications in cryptography. Here, given a sequence $s$ of $N$ integers, the goal is to find a duplicate in $s$. If one has the guarantee that $\Theta(N)$ elements of the sequence are duplicated, which is the case for instance when the sequence consists of random integers from $[N]$, it is well-known that classically $\Theta(\sqrt{N})$ queries are necessary and sufficient due to the birthday paradox. In the quantum model, this can be solved with query complexity $\Theta(N^{1/3})$ by the algorithm of Brassard, Høyer and Tapp [BHT]. The matching lower bound was first stated for a specific set of hard instances known as $2$-to-$1$ (i.e. each integer appears exactly twice or not at all) by Aaronson and Shi [AaronsonShi]. For some constant integer $k\geq 3$, those results can be further extended for finding $k$-collisions in a random input with suitable alphabet size, so that it contains $\Theta(N)$ $k$-duplicates with high probability. The classical query complexity for this problem is $\Theta(N^{1-1/k})$ [hellweg2013property, peng2023optimal], and quantumly it is $\Theta\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$ [zhandry_collision]. The situation is more complex for non-random inputs.
Remarkably, the complexity of testing $k$-collision-freeness (i.e., the absence of $k$-collisions) is harder to settle on the lower bound side. In this work, our lower bound shows the hardness of distinguishing inputs that have linearly many collisions from those that do not have any. For $k=2$, the two problems have the same complexity, since intuitively the only way to distinguish is to find a collision. This can be formalized easily in the classical case. Quantumly, this is more challenging, but the lower bound in [AaronsonShi] proved the hardness of distinguishing between $2$-to-$1$ instances and ones with no duplicate.
However, for larger $k$, distinguishing such inputs might be easier than finding a collision. Classically the upper bound of $O(N^{1-1/k})$ is straightforward for the finding variant. In the lower bounds of [hellweg2013property, peng2023optimal], they consider the distinguishing version, so classically the question is settled. But in the quantum setting, the upper and lower bounds of [zhandry_collision] are tight only for finding $k$-collisions, and for the distinguishing variant, we are not currently aware of anything better than the $\Omega(N^{1/3})$ lower bound corresponding to the $k=2$ case. To our knowledge, this problem has not yet been studied in the quantum setting. Here, we present an almost matching lower bound.
Theorem 1.4 (Restated in Theorem 4.1).
@@ 1.5 Open problems
First, there is still a gap between our lower and upper bound on the quantum query complexity of testing $k$-collision-freeness. In [improved_kDist], the authors keep using the dual polynomial method to improve the lower bound of [polynomial_strikes_back] for the $k$-distinctness problem. They achieve this by using a slightly different dual polynomial for $\mathrm{THR}_{N}^{k}$, where they allow more weight on the false positive inputs. This makes it impossible to prove the high correlation of the dual and the primal function, so they use a modified block composition. Our technique might be combined with this other approach to improve our lower bound to $\tilde{\Omega}(N^{1/2-1/(4k)})$.
@@
In [czumaj2016relating] it was proved that if a graph property can be tested with $O(1)$ queries in the bidirectional model, then it can be tested using $O(N^{1-\Omega(1)})$ queries in the unidirectional model. It would be very interesting to investigate if it also implies a quantum tester with query complexity say $O(N^{1/2-\Omega(1)})$.
@@ 2.1 Notations and basic definitions
Let us denote $[n]=\{1,\dots,n\}$ and $[n]_{0}=\{0,\dots,n\}$. When dealing with Boolean variables, we will usually use $b\in\{-1,1\}$ instead of $b\in\{0,1\}$. We can get to one from the other easily with the mapping $b\mapsto 1-2b$, or its inverse, which means that $-1$ is going to be treated as the “true” or “accepting” value. The reason for using $\{-1,1\}$ is that when dealing with dual polynomials it is easier to use this notation.
@@
A directed graph or digraph $G=(V,E)$ is a pair of a vertex set $V$ and an edge set $E$. The latter consists of directed edges that are ordered pairs of vertices: we say that $(u,v)\in E$ is directed from $u$ to $v$ where $u,v\in V$. We say that there is a directed path from $s=v_{0}$ to $t=v_{l+1}$ (with $s,t\in V$) if there exists an integer $l$ and vertices $v_{1},\dots,v_{l}\in V$ such that $\forall i\in[l]_{0}:\ (v_{i},v_{i+1})\in E$. A digraph $G=(V,E)$ is called strongly connected if for every $u\in V$ and $v\in V\setminus\{u\}$, there exists a directed path from $u$ to $v$. A subgraph of a graph $G=(V,E)$ is any graph $G^{\prime}=(V^{\prime},E^{\prime})$ satisfying $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ and $E^{\prime}\subseteq V^{\prime}\times V^{\prime}$.
Finally, throughout this work, notations $O(\cdot)$ and $\Omega(\cdot)$ will be hiding the dependencies on parameters $\varepsilon$, $k$ and $d$ that we consider to be constants.
@@ 2.2 Query complexity
Query complexity measures the minimum number of queries that an algorithm has to make in order to decide whether a property $P:D_{P}\subseteq\Sigma^{I}\to\{-1,1\}$ is satisfied, for an arbitrary input $x\in D_{P}$. Since the work of [beals_poly], it has been known that in the Boolean case (i.e. when $\Sigma=\{-1,1\}$), the acceptance probability of a $T$-query bounded-error quantum algorithm is a multivariate polynomial (in all the $x_{i}$’s) of degree at most $2T$.
@@
For completeness, we note that in some of the previous work on the unidirectional model they do impose the degree bound on both the out- and in-degree [czumaj2016relating]. This is mostly because this makes for an easier comparison between the uni- and bidirectional models, as this way they allow the same set of graphs. In this work we only assume that only the out-degrees are bounded by $d$.

### v2

@@ 1.4 Our results
1.5 Technical overview
1.5.1 Subgraph-finding algorithm
We start by describing how to prove the upper bound result of Theorem 1.1 for testing $k$-source-subgraph-freeness. We view the problem as a generalization of the problem of finding $k$-collisions and adapt an existing quantum algorithm for the latter problem. In [zhandry_collision], an algorithm is given for finding $k$-collisions in length-$N$ sequences of integers that contain $\Omega(N)$ $k$-collisions (e.g. $k$-to-$1$, or random sequence with appropriate parameters). Their algorithm generalizes the well-known collision finding algorithm of [BHT]. On a high level, the [zhandry_collision] algorithm first finds several $2$-collisions using Grover search like in [BHT], extends some of them to $3$-collisions in a similar way, and so on until a $k$-collision is found.
On the one hand, instead of random inputs, we consider the problem in the property testing context; and on the other hand, we generalize collision-finding to subgraph-finding. As a first step let us look at what happens when we consider the property testing version of the $k$-collision problem. In order to be able to use the algorithm of [zhandry_collision], we have to prove that if a length-$N$ sequence is far from $k$-collision-freeness then it contains many $k$-collisions. Notice that the collisions are not necessarily distinct: if the input only contains the same integer $N$ times, it only contains one huge collision, but it is still $\varepsilon$-far from $k$-collision-freeness for any $\varepsilon<1-k/N$. Thus, what we need to show is that there are $\Omega(N)$ many disjoint size-$k$ sets of indices such that for each set, the sequence contains the same character at the positions of the set. This statement is true because otherwise, by modifying all the characters that are in positions contained in a set ($o(N)$ characters in total), we could get a $k$-collision-free sequence which contradicts being far from $k$-collision-freeness.
When we make the second step of turning to testing of subgraph-freeness, we need to prove a variant of this statement: if an $N$-vertex graph $G$ is far from $H$-freeness (for some constant-sized subgraph $H$) then it contains $\Omega(N)$ many “source-disjoint” $H$-subgraphs. This means that there are $\Omega(N)$ many such $H$-subgraphs in $G$ that the set of vertices in the source components of each $H$-subgraph are disjoint. We prove this fact in Proposition 3.2 and this allows us to further generalize the approach of [zhandry_collision]: first find several partial solutions where only a few source components of an $H$-subgraph are explored, and gradually extend these (using Grover search coupled with constant-depth BFS) until a complete $H$-subgraph is found.
Notice that this way our algorithm finds an $H$-subgraph in $G$ promised that $G$ is far from $H$-freeness. This task is at least as difficult as property testing, where the algorithm only has to distinguish whether $G$ is $H$-free or far from any $H$-free graph. So our algorithm provides an upper bound on the property testing variant of $H$-freeness.
1.5.2 Collision-freeness lower bound
Now we will discuss our approach to proving the lower bounds of collision-freeness (Theorem 1.2) and $k$-source-subgraph-freeness (Theorem 1.3). We first give a simple reduction from $k$-collision-freeness to $k$-star-freeness, which is a special case of $k$-source-subgraph-freeness. This way, it is enough to prove a lower bound on testing $k$-collision-freeness, and it implies the same result on testing $k$-source-subgraph-freeness. Since our lower bound approach crucially depends on the (dual) polynomial method, let us start by briefly discussing it.
The (dual) polynomial method
The polynomial method is a common way to prove quantum query complexity lower bounds. It relies on the fact that the acceptance probability of a $T$-query bounded-error quantum algorithm is a polynomial of degree at most $2T$ [beals_poly]. This way, for proving a quantum query complexity lower bound on calculating a function $f$, it suffices to argue that any approximating polynomial of $f$ has large degree. One of the key properties that such lower bounds exploit is the symmetry that the function $f$ may exhibit, such as invariance under some permutation of the input. For example, the first tight lower bound of $\Omega(n^{1/3})$ for the collision problem was proved in this way [AaronsonShi].
The polynomial method can be written in the form of a linear program, of which one can take the dual. By weak LP-duality, when using this dual characterization for proving a lower bound on function $f$, one needs to provide a “witness” of the approximating polynomial’s high degree, say $\Delta$. This witness is called the dual polynomial $\psi$ and, in the easiest case of total Boolean [^1] We use $\{-1,1\}$ where $-1$ corresponds to the “true value”. functions $f:\{-1,1\}^{n}\to\{-1,1\}$, it needs to have three properties:
  1. (i)
    High correlation with $f$: $\sum_{x}f(x)\psi(x)>\delta$;
  2. (ii)
    Normalization: $\sum_{x}|\psi(x)|=1$;
  3. (iii)
    Pure high degree $\Delta$: $\sum_{x}p(x)\psi(x)=0$, for every polynomial $p$ with degree $<\Delta$,
where the summations are all over ${x\in\{-1,1\}^{n}}$.
When the function $f$ is partial, i.e., only defined on a subset $D\subset\{-1,1\}^{n}$, there is some subtlety that could be handled in two ways (or even in a mixture or both): zero-out the dual polynomial outside $D$ (corresponding to “unbounded degree”); or rewrite condition (i) accordingly (corresponding to “bounded degree”):
  1. (i’)
    High correlation with $f$: $\sum_{x\in D}f(x)\psi(x)-\sum_{x\not\in D}|\psi(x)|>\delta$.
Collision function
The paper of [polynomial_strikes_back] also used the dual polynomial method for proving quantum lower bounds for many problems, most of them being open before that work. Similarly to that paper, we need to take several steps to be able to use the dual polynomial method for the problem of property testing $k$-collision-freeness. This problem was not addressed in [polynomial_strikes_back].
One of the main conceptual ideas in [polynomial_strikes_back] is to re-formulate the problem we study as a composition of two simple Boolean functions. In that paper, powerful techniques are also developed in order to design dual polynomials for simple functions that can be composed. A common way of composing dual polynomials (called dual block composition) dates back to [ShiZ09, Lee09, Sher13], but [polynomial_strikes_back] provides new tools for handling it efficiently. We are going to reuse some of them, and also extend one in a way.
The first step is to find the right problem that can fit in the framework. We introduce a partial symmetric function $F$ defined on input strings $s=(s_{1},\dots,s_{N})\in[R]^{N}$. The domain of $F$ corresponds to the following promise: either $F$ has no $k$-collision, or it has many $k$-collisions occurring for distinct values. More formally,
$F(s)=\begin{cases}-1&\text{if no integer occurs at least }k\text{ times in }s,\\ 1&\text{if more than }\gamma R\text{ distinct integers occur at least }k\text{times in }s,\\ \text{undefined}&\text{otherwise}.\end{cases}$
This partial function is not a property testing problem, however it corresponds to a special case of testing $k$-collision-freeness, which is therefore enough to prove lower bounds.
Binary encoding
Now we encode the input string $s=(s_{1},\dots,s_{N})\in[R]^{N}$ into binary variables $x_{i,j}$ storing whether $s_{i}=j$, as in [Aaronson02]. Doing so, starting from the function $F$ above, we end up with a function $f$ defined over binary variables satisfying several symmetries, under the permutation of either $i$ or $j$ in $x_{i,j}$.
Moreover, the symmetries of $f$ allow the extension of the initial function $f$ from the very restricted set of binary inputs correspond to valid strings, to the more general set of binary inputs of Hamming weight $N$ [ambainis2016efficient]. With further technicalities one can also extend $f$ to all binary inputs of Hamming weight at most $N$ [BT20]. This is fundamental because instead of being forced to zero out the dual polynomial outside the domain of $f$, we only need to do so on inputs of Hamming weight higher than $N$. Using the symmetry of $f$, it can be shown that a lower bound on this modified, Boolean version implies a lower bound on the original $k$-collision problem.
This way we end with two promises on the binary encoding of the input. The first one comes from the function $F$ itself: we have the promise that the input contains either no $k$-collision or it has many ones with different values. The second promise is the consequence of the encoding: we want the binary encoding to have Hamming weight at most $N$. Let $D$ denote the set of binary strings satisfying both promises, and let $H_{\leq N}$ denote the set of binary strings with Hamming weight at most $N$. For this case we use the “double-promise” version of the dual polynomial method, where, in order to prove that every $\delta$-approximating polynomial of $f$ has degree at least $\Delta$, the dual polynomial has to satisfy four conditions, where the fourth one corresponds to zeroing out $\psi$ on large Hamming weight inputs [polynomial_strikes_back]:
  1. (i’)
    High correlation with $f$: $\sum_{x\in D}f(x)\psi(x)-\sum_{x\in H_{\leq N}\setminus D}|\psi(x)|>\delta$;
  2. (ii)
    Normalization: $\sum_{x}|\psi(x)|=1$;
  3. (iii)
    Pure high degree $\Delta$: $\sum_{x}p(x)\psi(x)=0$, for every polynomial $p$ with degree $<\Delta$;
  4. (iv)
    No support on inputs with large Hamming weight: $\psi(x)=0$, for every $x\notin H_{\leq N}$.
Composition
Coming back now to the definition of our Boolean function $f$, one can rewrite it as a composition of simpler functions: $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$, where by composition we mean $(g\circ h)(x)=g(h(x_{1}),\dots,h(x_{n}))$ (where $x=(x_{1},\dots,x_{n})$ and each $x_{i}$ is a binary vector of appropriate dimension) and the domain is restricted to bit strings of Hamming weight at most $N$. Note that here, $x_{j}=(x_{1,j},x_{2,j},\ldots,x_{N,j})$. Above, $\mathrm{THR}_{N}^{k}$ is the threshold function: it is $-1$ if the input bitstring contains at least $k$ many $-1$ (true) values, and is $1$ otherwise; and $\mathrm{GapOR}_{R}^{\gamma}$ is the gap version of $\mathrm{OR}$, which is $1$ if the input is all-1, $-1$ if the input contains at least $\gamma R$ many $-1$ values, and is undefined otherwise.
In order to give a dual polynomial for this composed function, we start from a dual polynomial for each part of the composition ($\phi$ and $\psi$), which were already given in [polynomial_strikes_back] (in different contexts). Then we use a known way [ShiZ09, Lee09, Sher13] of composing dual polynomials called the dual block composition, which provides a nearly good dual polynomial $\phi\star\psi$ for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$. Indeed, by construction, the normalization (ii) and the pure high degree (iii) are guaranteed. However, the issue is that it is not 00 on bitstrings of large Hamming weight thus (iv) is not satisfied, and and the high correlation (i’) still has to be proved.
To fix (iv), we use another result of [polynomial_strikes_back] which provides another dual polynomial $\zeta$, that is close to $\phi\star\psi$ and that is 00 on inputs having Hamming weight larger than $N$. Also, it only changes the pure high degree by a polylogarithmic factor. Now the only remaining task is to prove a large enough correlation (i’) of $\phi\star\psi$ and $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ so that $\zeta$ still has high enough correlation. This high correlation proof (Lemma 4.18) is the most technical part of our paper.
High correlation: proof of Lemma 4.18
The statement we prove is the following high correlation bound:
$\sum_{x\in D}(\phi\star\psi)(x)\cdot(\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k})(x)-\sum_{x\notin D}|(\phi\star\psi)(x)|\geq 9/10.$
In the proof of this lemma we use Proposition 4.17 that is a more general statement of some techniques used in several proofs of [polynomial_strikes_back]. But then we need to diverge from their proof because it crucially relies on a certain one-sided error property (in the sense of [polynomial_strikes_back, Lemma 6.11]) of the inner function of the composition which is OR function in their case. Our inner function is the threshold function which does not satisfy this property, so we have to use some other properties of the dual polynomials in our proof. This different proof technique could be a step towards obtaining a more general lower bound technique.
In fact, the authors in [ambainis2016efficient] stated it as an open question if one could use a variant of the proportional moments technique for proving better quantum lower bounds. We leave this question open, and conjecture that a similar result holds in the quantum setting with a lower bound of $\Omega\Big{(}N^{\frac{1}{2}\left(1-\frac{1}{k}\right)}\Big{)}$. One can consider this work as a proof of this conjecture for the special case of $k$-collision-freeness, and we hope that it will serve as a step towards proving it in general.
Comparison with [polynomial_strikes_back]
The structure and several elements of our lower bound proof come from [polynomial_strikes_back]. Using their work is it relatively easy to consider the binary encoding of the $k$-collision function, extend its domain and relate it to the composition of easier functions. They also provide dual polynomial we can use and we can compose them using the dual block composition. We can even use another result of theirs that provides a dual polynomial that is zero on large Hamming weight inputs. Then the only remaining task is to prove the high correlation, for which we need to diverge from this paper because their respective proof relies on a one-sided error property that does not hold for our problem. This way we prove our result in a more difficult, two-sided error setting.
1.5.3 3-colorability lower bound
Let us now discuss our approach to proving the linear lower bound on quantum query complexity of testing $3$-colorability (Theorem 1.4). Before proceeding to present our approach, let us briefly discuss the classical lower bound of testing $3$-colorability. To prove the classical lower bound of $3$-colorability, the authors in [bogdanov2002lower] first studied another problem called E$(3,c)$LIN-2, a problem related to deciding the satisfiability of a system of linear equations. More formally, E$(3,c)$LIN-2 considers a system of linear equations modulo $2$, where each equation has $3$ variables and every variable appears in at most $c$ equations. Given such a system of linear equations, the goal is to distinguish if it is satisfiable, or at least some suitable fraction of the equations need to be modified to satisfy it. [bogdanov2002lower] proved that $\Omega(N)$ classical queries to the system of linear equations are necessary for testing E$(3,c)$LIN-2.
After this, they designed a reduction from E$(3,c)$LIN-2 to $3$-colorability such that satisfying instances of E$(3,c)$LIN-2 are reduced to $3$-colorable graphs, and far from satisfiable instances of E$(3,c)$LIN-2 are mapped to far from $3$-colorable graphs. Combining these two arguments, the authors in [bogdanov2002lower] proved that $\Omega(N)$ classical queries are necessary for testing $3$-colorability for bounded degree graphs.
The authors in [bogdanov2002lower] used Yao’s minimax method to prove the linear lower bound in testing E$(3,c)$LIN-2. In particular, they designed two distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ such that the systems of linear equations in $D_{\mathrm{yes}}$ are satisfiable, whereas the systems of linear equations in $D_{\mathrm{no}}$ are far from being satisfiable. A crucial ingredient of their lower bound proof is a construction of a system of linear equations (represented as a matrix) that are far from being satisfiable, but any $\delta N$ rows of the matrix are linearly independent. Hence, any subset of $\delta N$ entries of the matrix will look uniformly random, and therefore hard to distinguish from a satisfiable instance.
It is a known fact that distinguishing between a uniformly random string and a $\ell$-wise independent string, for an appropriate integer $\ell$, is hard for quantum algorithms (see e.g. [apers2022quantum]). Using this result, we can construct suitable hard instances for E$(3,c)$LIN-2 such that testing E$(3,c)$LIN-2 remains maximally hard (requires $\Omega(N)$ queries) for any quantum algorithm. Combining this hardness result with the reduction from E$(3,c)$LIN-2 to $3$-colorability, we finally prove that $\Omega(N)$ quantum queries are necessary for testing $3$-colorability. We formally prove this in Section 5.
@@ 1.5.2 Collision-freeness lower bound
Later, in [yoshida2010query], the authors used various reductions to $3$-colorability to argue that a number of other important problems including testing Hamiltonian Path/Cycle, approximating Independent Set/Vertex Cover size etc, are maximally hard to test in the classical model. As a corollary of our quantum lower bound, we also obtain maximal quantum query complexity for these problems.
@@ 1.6 Open problems
Our work raises several important open questions. First, there is still a gap between our lower and upper bounds on the quantum query complexity of testing $k$-collision-freeness. In [improved_kDist], the authors keep using the dual polynomial method to improve the lower bound of [polynomial_strikes_back] for the $k$-distinctness problem. They achieve this by using a slightly different dual polynomial for $\mathrm{THR}_{N}^{k}$, where they allow more weight on the false positive inputs. This makes it impossible to prove the high correlation of the dual and the primal functions, so they use a modified block composition. Our technique might be combined with this other approach to improve our lower bound to $\tilde{\Omega}(N^{1/2-1/(4k)})$.
@@
In [czumaj2016relating], it was proved that if a graph property can be tested with $O(1)$ queries in the bidirectional model, then it can be tested using $O(N^{1-\Omega(1)})$ queries in the unidirectional model. It would be very interesting to investigate if it also implies a quantum tester with query complexity, say $O(N^{1/2-\Omega(1)})$.
@@ 2.1 Notations and basic definitions
Let us denote $[n]=\{1,\dots,n\}$ and $[n]_{0}=\{0,\dots,n\}$. When dealing with Boolean variables, we will usually use $b\in\{-1,1\}$ instead of $b\in\{0,1\}$. We can get to one from the other easily with the mapping $b\mapsto 1-2b$, or its inverse, which means that $-1$ is going to be treated as the “true” or “accepting” value. The reason for using $\{-1,1\}$ is that when dealing with dual polynomials, it is easier to use this notation.
@@
A directed graph or digraph $G=(V,E)$ is a pair of a vertex set $V$ and an edge set $E$. The latter consists of directed edges that are ordered pairs of vertices: we say that $(u,v)\in E$ is directed from $u$ to $v$ where $u,v\in V$. We say that there is a directed path from $s=v_{0}$ to $t=v_{l+1}$ (with $s,t\in V$) if there exists an integer $\ell$ and vertices $v_{1},\dots,v_{\ell}\in V$ such that $\forall i\in[\ell]_{0}:\ (v_{i},v_{i+1})\in E$. A digraph $G=(V,E)$ is called strongly connected if for every $u\in V$ and $v\in V\setminus\{u\}$, there exists a directed path from $u$ to $v$. A subgraph of a graph $G=(V,E)$ is any graph $G^{\prime}=(V^{\prime},E^{\prime})$ satisfying $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ and $E^{\prime}\subseteq V^{\prime}\times V^{\prime}$.
Finally, throughout this work, notations $O(\cdot)$ and $\Omega(\cdot)$ will be hiding the dependencies on parameters $\varepsilon$, $k$ and $d$ that we consider to be constants. Additionally, we will use $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$, where we hide poly-logarithmic dependencies on the parameters.
@@ 2.2 Query complexity
Query complexity measures the minimum number of queries that an algorithm has to make in order to decide whether a property $\mathcal{P}:D_{\mathcal{P}}\subseteq\Sigma^{I}\to\{-1,1\}$ is satisfied, for an arbitrary input $x\in D_{\mathcal{P}}$. Since the work of [beals_poly], it has been known that in the Boolean case (i.e., when $\Sigma=\{-1,1\}$), the acceptance probability of a $T$-query bounded-error quantum algorithm is a multivariate polynomial (in all the $x_{i}$’s) of degree at most $2T$.
@@
For completeness, we note that in some of the previous works on the unidirectional model, the authors do impose the degree bound on both the out- and in-degree (see, e.g., [czumaj2016relating]). This is mostly because this makes for an easier comparison between the uni- and bidirectional models, as this way they allow the same set of graphs. In this work, we assume that only the out-degrees of the vertices are bounded by $d$.